OpenStudy (anonymous):

A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = ; 1 mile = 1609 meters.

OpenStudy (anonymous):

this is kinda hard

OpenStudy (anonymous):

70/3 to get miles per hour. But it does not say if it should be meters per second miles per minute inches per day... what exactly is the speed unit conversion you need?

OpenStudy (anonymous):

A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = ; 1 mile = 1609 meters

OpenStudy (anonymous):

thats all it said

OpenStudy (anonymous):

these are answers u can choose from

OpenStudy (anonymous):

10.4 meters per second 23.3 meters per second 52.2 meters per second 93.9 meters per second

OpenStudy (anonymous):

ok 70mi * 1609m * 1 1 hr 1 min ------- ----- * ----- * ----- 1 mi 3hr 60min 60 sec miles cancel out from top and bottom hr cancel out from top and bottom min cancel out from top and bottom leaving meters per second multiply straight across to find answer

OpenStudy (anonymous):

obviously you may need to reduce

OpenStudy (anonymous):

ill try but can u try to find the answer please

OpenStudy (anonymous):

i have found it (70*1609) --------- = ? (3*60*60)

OpenStudy (anonymous):

I see the answer amoung those you liseted

OpenStudy (anonymous):

do you have a calculator handy?

OpenStudy (anonymous):

no i dont have a caculator

OpenStudy (anonymous):

you know there is one built into windows...

OpenStudy (anonymous):

IHAVE A CACULATOR SO WAT AM I SUPPOST TO DO

OpenStudy (anonymous):

(70*1609) --------- = ? (3*60*60)

OpenStudy (anonymous):

70*1609 = x 3*60*60 = y x/y = your needed answer

OpenStudy (anonymous):

what do you come up with?

OpenStudy (anonymous):

THIS IS FOR REVIEW PLEASE HELP I have an oral exam at 12 and need these answered help please!!! :0 How do you define the appropriate quantities to model a situation or description? How do you choose the level of accuracy given the limitations of a situation? THIS IS FOR REVIEW

OpenStudy (anonymous):

I DONT KNOW HOW TO DO IT ON THE CAULATOR HELP !